OpenStudy (anonymous):

ok I need help with this problem and its due tonight. its goes as such below: short leg= x long leg= x+7 Longest leg= 35 this is a squared plus b squared equals c squared and i am hopeless at math help please :/

OpenStudy (anonymous):

a² + b² = c² x² + ( x +7)² = 35²

OpenStudy (anonymous):

Ok. Where do I go from there? Apologies math is not my best area.

OpenStudy (anonymous):

Expand ( x +7)² = ....

OpenStudy (anonymous):

(x+7)(x+7)?

OpenStudy (anonymous):

FOIL them out, so that you can simplify with other terms!

OpenStudy (anonymous):

Ok that is how far I got then I got lost. 2x^2+14x+49=1225

OpenStudy (anonymous):

Great job, indeed :)

OpenStudy (anonymous):

but i need the shortest leg

OpenStudy (anonymous):

You need the form ax² + bx + c = 0 to solve for x :)

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

I thought we were using the Pythagoreom Theorom.

OpenStudy (anonymous):

It is college math. I barely passed in high school.

OpenStudy (anonymous):

We did at the first step :)

OpenStudy (anonymous):

I'm really sorry I get mixed up easily :C

OpenStudy (anonymous):

2x² + 14x + 49 = 1225 Can you make the right side "0"

OpenStudy (anonymous):

that would end up with the single number being negative is that ok?

OpenStudy (anonymous):

er well if it is allowed to have a negative in there then it would be......

OpenStudy (anonymous):

negative 1176

OpenStudy (anonymous):

which would make the equation 2x^2+14x-1176

OpenStudy (anonymous):

2x² + 14x -1176 = 0 Can you solve for x now?

OpenStudy (anonymous):

i think so let me try apologies again aspbergers does not make abstract concepts easy in the least

OpenStudy (anonymous):

28,21

OpenStudy (anonymous):

x = 21!

OpenStudy (anonymous):

ok let me try that and see if it works

OpenStudy (anonymous):

awesome thank you thank you

OpenStudy (anonymous):

Omit the negative value x = -28 ( doesn't make sense for the length )

OpenStudy (anonymous):

it was positive 28 but 21 was right thank you for your help

OpenStudy (anonymous):

No, x = -28, not 28 !

OpenStudy (anonymous):

28 worked as well as 21

OpenStudy (anonymous):

thank you again for your help :)

OpenStudy (anonymous):

ahhh well thank you again. have a good day :)